40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 9 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  16462   Tue Nov 9 18:05:03 2021 Ian MacMillanSummaryComputer Scripts / ProgramsSUS Plant Plan for New Optics

[Ian, Tega]

Now that the computer is in its new rack I have copied over the filter two files that I will use in the plant and the controller from pianosa:/opt/rtcds/caltech/c1/chans to the docker system in c1sim:/home/controls/docker-cymac/chans. That is to say, C1SUP.txt -> X1SUP.txt and C1SUS.txt -> X1SUS_CP.txt, where we have updated the names of the plant and controller inside the txt files to match our testing system, e.g. ITMX -> OPT_PLANT in plant model and ITMX -> OPT_CTRL in the controller and the remaining optics (BS, ITMY, PRM, SRM) are stripped out of C1SUS.txt in order to make X1SUS_CP.txt. 

Once the filter files were copied over need to add them to the filters that are in my models to do this I run the commands:

$  cd docker-cymac
$  eval $(./env_cymac)
$  ./login_cymac
 #  cd /opt/rtcds/tst/x1/medm/x1sus_cp
 #  medm -x X1SUS_OPT_PLANT_TM_RESP.adl

see this post for more detail

Unfortunately, the graphics forwarding from the docker is not working and is giving the errors:

arg: X1SUS_OPT_PLANT_TM_RESP.adl

locateResource 'X1SUS_OPT_PLANT_TM_RESP.adl'

isNetworkRequest X1SUS_OPT_PLANT_TM_RESP.adl

canAccess('X1SUS_OPT_PLANT_TM_RESP.adl', 4) = 0

can directly access 'X1SUS_OPT_PLANT_TM_RESP.adl'

isNetworkRequest X1SUS_OPT_PLANT_TM_RESP.adl

locateResource(X1SUS_OPT_PLANT_TM_RESP.adl...) returning 1

Error: Can't open display:

This means that the easiest way to add the filters to the model is through the GUI that can be opened through X2go client. It is probably easiest to get that working. graphics forwarding from inside the docker is most likely very hard. 

unfortunately again x2go client won't connect even with updated IP and routing. It gives me the error: unable to execute: startkde. Going into the files on c1sim:/usr/bin and trying to start startkde by myself also did not work, telling me that there was no such thing even though it was right in front of me.

  16461   Tue Nov 9 16:55:52 2021 Ian MacMillanSummaryComputer Scripts / ProgramsSUS Plant Plan for New Optics

[Ian, Tega]

We have moved c1sim computer from the test stand to the server rack in the office area. (see picture)

It is connected to the general campus network. Through the network switch at the top of the rack. This switch seeds the entire Martian network.

Test to show that I am not lying:

  1. you can ping it or ssh into it at
    controls@131.215.114.116
    Using the same password as before. Notice this is not going through the nodus network.
  2. It also has a different beginning of the IP addresses. Martian network IP addresses start with 191.168.113

c1sim is now as connected to the 40m network as my mom's 10-year-old laptop.

unfortunately, I have not been able to get the x2go client to connect to it. I will have to investigate further. It is nice to have access to the GUI of c1sim occasionally.

Attachment 1: IMG_8107.JPG
IMG_8107.JPG
  16460   Tue Nov 9 13:40:02 2021 Ian MacMillanSummaryComputer Scripts / ProgramsSUS Plant Plan for New Optics

[Ian, Tega]

After talking with Rana we have an updated plan. We will be working on this plan step by step in this order.

  1. Remove c1sim from the test stand rack and move it to the rack in the office next to the printer. When connecting it we will NOT connect it to the Martian network! This is to make sure that nothing is connected to the 40m system and we can't mess anything up.
  2. Once we have moved the computer over physically, we will need to update anyone who uses it on how to connect to it. The way we connect to it will have changed.
  3. Now that we have the computer moved and everyone can connect to it we will work on the model. Currently, we have the empty models connected.
    1. recompile the model since we moved the computer.
    2. verify that nothing has changed in the move and the model can still operate and compile properly
  4. The model has the proper structure but we need to fill it with the proper filters and such
    1. For the Plant model
      1. To get it up and running quickly we will use the premade plant filters for the plant model. These filters were made for the c1sup.mdl and should work in our modified plant model. This will allow us to verify that everything is working. And allow us to run tests on the system.
      2. We need to update the model and add the state space block. (we are skipping this step for now because we are fast-tracking the testing)  
        1. Check with Chris to make sure that this is the right way to do it. I am pretty sure it is, but I don't know anything
        2. Make the 6 DOF state-space matrix. We only have a three DOF one. The surf never made a 6 DOF. 
        3. Make the block to input into the model
        4. make a switch that will allow us to switch between the state-space model and the filter block
    2. For the controller
      1. Load filter coefficients for the controller model from one of the current optics and use this as a starting point.
      2. Add medm screens for the controller and plant. We are skipping this for now because we want results and we don't care if the screens look nice and are useable at the moment.
  5.  Test the model
    1. we will take an open-loop transfer function of all six of the DOFs to all other DOFs which will leave us with 36 TFs. Many will be zero
      1. If you are looking at this post then we are measuring transfer functions from the blue flags to the green flags across the plant model.
      2. We will want to look at the TFs across the controller

 

  16459   Tue Nov 9 11:11:37 2021 YehonathanUpdateBHDSOS assembly

The gluing was mostly successful. Only two magnets didn't stick (see attachment).

Attachment 1: 20211109_110408.jpg
20211109_110408.jpg
  16458   Mon Nov 8 18:42:38 2021 KojiSummaryBHDRack Layout / Power Strips
Rack # of units that
requires +18V
Power Source
1X3 (new rack) 15 1X3 U1/2
1X4 13 1X3 U1/2
1X5 8 or 9 (OL AA) 1X5 U40/41
1Y0 17 1Y0 U1/2
1Y1 15 1Y0 U1/2
1Y3 12 1Y3 U39/40
1X9 9 1X9 U38/39
1Y4 9

1Y4 U39/40

Notes:

  • There are 8 racks and there is only 7x 18V power strips. 1X5 could be the one without the power strip and to parasite with 1X3/4. Otherwise we need to modify some of the 24V power strips (no plan to use) into 18V by replacing the connectors.
  • We need total ~100 18V cables / We ordered 60x 3ft / 60x 3ft / 30x 10ft. Hopefully these are enough for our depand... I haven't checked the delivered number.
  • All the acromags are supposed to be powered with one voltage. I think they are supposed to run with +18V.
  • I didn't check the distribution of Sorensens through the lab. (i.e. how many we have / how many we need / ...)
Attachment 1: rack_plan.pdf
rack_plan.pdf
  16457   Mon Nov 8 17:52:22 2021 Ian MacMillanUpdateSUSSetting up suspension test model

[Ian, Tega]

We combined a controler and a plant model into a single modle (See first attachment) called x1sus_cp.mdl in the userapps folder of the cymac in c1sim. This model combines 2 blocks: the controler block which is used to control the current optics and is found in cvs/cds/rtcds/userapps/release/sus/c1/models/c1sus.mdl further the control block we are using comes from the same path but from the c1sup.mdl model. This plant model is the bases for all of my custom plant models and thus is a good starting point for the testing. It is also ideal because I know it can beeasily altered into a my state-space plant model. However, we had to make a few adjustments to get the model up to date for the cds system. So it is now a unique block.

These two library blocks are set in the userapps/lib folder on the cymac. This is the lib file that the docker system looks to when it is compiling models. For a quick overview see this. All other models have been removed from the MatLab path so that when we open x1sus_cp.mdl in MatLab it is using the same models it will compile with.

We could not find the rtbitget library part, but chris pointed us to userapps, and we copied it over using: scp /opt/rtcds/userapps/trunk/cds/common/models/rtbitget.mdl controls@c1sim:/home/controls/simLink/lib.

NOTE TO FUTURE IAN: don't forget that unit delays exist.

Next step: now that we have a model that is compiling and familiar we need to make medm screens. We will use the auto mdl2adl for this so that it is quick. Then we can start adding our custom pieces one by one so that we know that they are working. We will also work with Raj to get an independent python model working. Which will allow us to compare the cds and python models.

Attachment 1: x1sus_cp.png
x1sus_cp.png
  16456   Mon Nov 8 17:22:27 2021 YehonathanUpdateBHDSOS assembly

Big Gluing Day

Today I glued the magnet+dumbell assemblies on the optics adapters.

Unlike magnet gluing on a 3" optic where one can use a magnet gluing fixture, here I had to position the magnets manually. There is a complication though: the magnet is much heavier than the dumbell making it almost impossible gluing the dumbell side down onto the adapter since it is very unstable in this position. A workaround is to put the magnets on some paramagnetic sheet so that the magnets stick to it and then flip it over and glue it on the adapter dumbell sides down.

The problem here is that I need to position the magnets relatively accurately on the metal sheet. To make things slightly easier I printed some drawings of the positions of the magnet, laminated them, and cleaned them to have a decent starting point (attachment 1).

For each adapter:

1. I applied glue to the 4 circular grooves at the back of the adapter.

2. I picked 4 magnets (2 north, 2 south). Trying to match their strength.

3. Made a note of which magnets I picked for which adapter in the magnet+dumbell spreadsheet.

4. Clean the dumbells' surfaces when necessary.

5. Put the magnets on a laminated magnet-positions-drawing on a metal sheet that was precleaned in the right order.

6. Flip the metal sheet and position it on the adapter such that the dumbells go as precisely as possible into the circular grooves on the adapater.

7. Adjust the magnets' positions by pushing them slightly with a non-magnetic tip.

Attachment 2 shows the numbering on the adapters for future tracking.

I also glued some magnets and aluminum rods to side blocks. Next gluing session I will glue magnets to the aluminum rods. Probably some dumbells will not stick well to the adapters. These will have to be cleaned and reglued as well.

 

 

Attachment 1: Magnets_Positions.png
Magnets_Positions.png
Attachment 2: Adapters_Order.png
Adapters_Order.png
  16455   Mon Nov 8 15:29:05 2021 PacoSummaryBHD1Y1 rack work; New power for cameras

[Paco, Anchal]

In reference to Koji's concern (see previous elog), we have completely removed sorensen power supplies from 1Y1. We added a 12 Volts / 2 Amps AC-to-DC power supply for the cameras and verified it works. We stripped off all unused hardware from shutters and other power lines in the strips, and saved the relays and fuses.

We then mounted SR2, PR3, PR2 Sat Amps, 1Y1 Sat amp adapter, and C1SUS2 AA (2) and AI (3) boards. We made all connections we could make with the cables from the test stand, as well as power connections to an 18 VDC power strip.

  16454   Mon Nov 8 13:13:00 2021 KojiSummaryBHD1Y1 rack work; Sorensens removed

Updated the rack layout. Now there is an issue.
We were supposed to have 1U space at the top, but it was occupied by the 12V.
We need to either lower the c1sus2 and IO chassis 1U or move the Sorensen at the bottom.

Attachment 1: 40m_BHD.png
40m_BHD.png
  16453   Mon Nov 8 10:13:52 2021 PacoSummaryBHD1Y1 rack work; Sorensens removed

[Paco, Chub]

Removed all sorensen power supplies from this rack except for 12 VDC one; that one got pushed to the top of the rack and is still powering the cameras.

  16452   Fri Nov 5 20:35:10 2021 KojiUpdateBHDFeedthru / Optical Mounts

- New feedthrus [4xDB25 Qty 4 / 8xDB25 Qty 1] are placed on the wire shelf at the entrance -> Jordan, please clean them.

- There are plenty of 2" DLC mounts. There are also many 1.5" mounts but they are less useful.
  We need at least 3 1" moounts and 1 1" or 2" lens mount (and the lens). Let's purchase them on Thorlabs. I'll work on the order.

Attachment 1: P_20211105_200817.jpg
P_20211105_200817.jpg
  16451   Fri Nov 5 12:49:32 2021 ranaUpdateSUSSetting up suspension test model

Please don't put it on c1sus2. Put it on the completely independent test stand as we discussed Wednesday. You must test the controller on the simplant and verify that they thing is stable and works, before putting it in the 40m network.

  16450   Fri Nov 5 12:21:16 2021 AnchalSummaryBHDPart VI of BHR upgrade - Removal of ITMYC optics

Today I opened the ITMY chamber and removed the following optics and placed them in Xend flow bench (See attachment 1-3 for updated photograph):

  • OM1
  • OM2
  • ITMYOL1
  • ITMYOL2
  • SRMOL1
  • SRMOL2
  • POYM1
  • 3 counterweights one of which was double the height of others.

I also unscrewed SRM and parked it near the Western end of the table where no optical paths would intersect it. Later we will move it in place once the alignment of the rest of the optics has been done.

While doing this work, I found two unnoted things on the table:

  • One mirror mounted on a mount but not on a post was just sitting next to ITMY. I have removed this and placed it on Xend flow bench.
  • One horizontal razor or plate on the South end of table, mounted on what I thought looked like a picomotor. The motor was soldered to wires without any connector in-line, so I could not remove this. This is on the spot of AS4 and will need to be removed later.

Photos: https://photos.app.goo.gl/S5siAYguBM4UnKuf8

Attachment 1: XendFlowBenchLeftEnd.jpg
XendFlowBenchLeftEnd.jpg
Attachment 2: XendFlowBenchMiddle.jpg
XendFlowBenchMiddle.jpg
Attachment 3: XendFlowBenchRightEnd.jpg
XendFlowBenchRightEnd.jpg
  16449   Thu Nov 4 18:29:51 2021 TegaUpdateSUSSetting up suspension test model

[Ian,Tega]

Today we continued working on setting up the 6 degrees of freedom model for testing the suspension which we copied over from  "/cvs/cds/rtcds/userapps/release/sus/c1/models/c1sup.mdl" to c1sp2.mdl in the same folder. We then changed the host from c1lsc to c1sus2, changed cpu # from 7 to 3 bcos c1sus2 has 6 cores. Then ran the following commands to build and install the model on c1sus2:

$ ssh c1sus2

$ rtcds make c1sp2

$ rtcds install c1sp2

where we encountered the following installation error:

ERROR: This node 62 is already installed as:

hostname=c1lsc

system=c1sup

The new entry you are trying to write is as follows:

hostname=c1sus2

system=c1sp2

This script will not overwrite existing entries in testpoint.par

If this is an attempt to move an existing system from one host to another,

please remove conflicting entry from testpoint.par file

It seems that changing the model name and host did not change the node allocation, so will remove the previous entries in testpoint.par to see if that helps. After deleting the following lines

[C-node62]
hostname=c1lsc
system=c1sup

from the file "/opt/rtcds/caltech/c1/target/gds/param/testpoint.par", the installation went fine and the above entries were replaced by 

[C-node62]
hostname=c1sus2
system=c1sp2

BTW, I now believe the reason we had the node conflict earlier was bcos both models still had the same value of  dcuid=62, so I think changing this value in our model file would be a better solution. Work is ongoing.

 

  16448   Thu Nov 4 15:03:43 2021 KojiSummaryBHD1Y1 rack work

I have visited the binder file for the 40m wiring file in the control room.
The 12V power supply on 1Y1 is for the CCD cameras. So we still want to keep the 12V 0.8A power and the side connections for these. It is not necessary to be Sorensen. Can we replace it with an AC-DC adapter with +12V/1A for example? BTW, the video matrix and quads are AC-powered.

The mysterious thick cables and cross-connects (green wires) on the side panel (labeled AP1/AP2/SP/IMCREFL) are for "EO shutters". It was meant for the protection of the PDs from bright beams.
I don't think they have been used. And we don't need them.

  16447   Wed Nov 3 16:55:23 2021 Ian MacMillanSummarySUSSUS Plant Plan for New Optics

[Ian, Tega, Raj]

This is the rough plan for the testing of the new suspension models with the created plant model. We will test the suspensions on the plant model before we implement them into the full

  • Get State-space matrices from the surf model for the SOS. Set up simplant model on teststand
    • The state-space model is only 3 degrees of freedom. (even the surf's model)
    • There are filter modules that have the 6 degrees of freedom for the suspensions. We will use these instead. I have implemented them in the same suspension model that would hold the state space model. If we ever get the state space matrices then we can easily substitute them.
  • Load new controller model onto test stand. This new controller will be a copy of an existing suspension controller.
  • Hook up controller to simplant.  These should form a closed loop where the outputs from the controller go into the plant inputs and the plant outputs go to the controller inputs.
  • Do tests on set up.
    • Look at the step response for each degree of freedom. Then compare them to the results from an existing optic. 
      • Also, working with Raj let him do the same model in python then compare the two.
  • Make sure that the data is being written to the local frame handler.

MEDM file location

/opt/rtcds/userapps/release/sus/common/medm/hsss_tega_gautam

run using 

For ITMX display, use:

hsss_tega_gautam>medm -x -macro "site=caltech,ifo=c1,IFO=C1,OPTIC=ITMX,SUSTYPE=IM,DCU_ID=21,FEC=45" SUS_CUST_HSSS_OVERVIEW.adl

  16446   Tue Nov 2 22:50:30 2021 KojiUpdateBHDSOS assembly

2.5Hz is surprising. Can you move it down to sub 1Hz by adding a socket cap screw at the top instead of the set screw for the Teflon piece? How much mass do you need to add?

  16445   Tue Nov 2 18:54:41 2021 Anchal, PacoUpdateBHDOptical fibres laid for BHD upgrade

We successfully laid down all required optical fibre fiber cables from 1X4-1X7 region to 1Y1-1Y3 region today. This includes following cables:

  • Timing fibre fiber from Master Timing Synchornizer D050239 on 1X6 to C1SU2 I/O chassis on 1Y1.
  • Timing fibre fiber from Master Timing Synchornizer D050239 on 1X6 to C1BHD I/O chassis on 1Y3.
  • CX4 cable from Dolphin Card on 1X4 to C1SU2 FE on 1Y1 for IPC.
  • CX4 cable from Dolphin Card on 1X4 to C1BHD FE on 1Y3 for IPC.
  • DAQ Network extension fibre fiber optic cable from DAQ Network Switch on 1X7 to another switch we mounted on 1Y3 for local DAQ network distribution.
  16444   Tue Nov 2 16:42:00 2021 PacoSummaryBHD1Y1 rack work

[paco, ian]

After the new 1Y0 rack was placed near the 1Y1 rack by Chub and Anchal, today we worked on the 1Y1 rack. We removed some rails from spaces ~ 25 - 30. We then drilled a pair of ~ 10-32 thru-holes on some L-shaped bars to help support the c1sus2 machine weight. The hole spacing was set to 60 cm; this number is not a constant across all racks. Then, we mounted c1sus2. While doing this, Paco's knee clicked some of the video MUX box buttons (29 and 8 at least). We then opened the rack's side door to investigate the DC power strips on it before removing stuff. We did power off the DC33 supplies on there. No connections were made to allow us to keep building this rack.

When coming back to the control room, we noticed 3/4 video feed (analog) for the Test masses had gone down... why?


Next steps:

  • Remove sorensen (x5) power supplies from top of 1Y1 .. what are they actually powering???
  • Make more bars to support heavy IO exp and acromag chassis.
  • Make all connections (neat).

Update Tue Nov 2 18:52:39 2021

  • After turning Sorensens back up, the ETM/ITM video feed was restored. I will need to hunt the power lines carefully before removing these.
  16443   Tue Nov 2 15:02:03 2021 YehonathanUpdateBHDSOS assembly

{Yehonathan, Tega}

We took the free-swinging spectra of the OpLev in the X and Y direction.

To make the motion of the optic quiet we turned off the airflow on the optical bench and moved the QPD close to the SOS so that the laser beam stays more or less within the QPD sensitive area.

In the process, we realized that the cleanroom HeNe went bad. It turned off after a few minutes after turning it on. The behavior repeated with another power supply. We replaced the HeNe and realigned it coarsely.

The data was taken using an oscilloscope while the optic was swinging freely. The PSD was calculated afterward (attachment 1).

Surprisingly, the pitch has a resonance frequency of ~ 2.5 Hz. And this is after we removed the back counterweight.

Additionally, we aligned the tilt of the optical table. Using a spirit bubble we adjusted the tilt by using a wrench on the table legs. As we suspected, the table was slightly tilted in the north-south direction.

Attachment 1: FreeSwingingSpectra.pdf
FreeSwingingSpectra.pdf
  16442   Mon Nov 1 14:51:34 2021 KojiUpdateGeneralChecking the vent plan

The vent team described a detailed vent plan (and reports where the actions have been performed)

https://wiki-40m.ligo.caltech.edu/vent/Fall2021

- [Sec.4] We should decide the final PR2 mirror through table-top measurements.

- [Sec 6] BS alignment is probably "unknown" now. So it'd be better to use the ITMY spot as the reference, then align BS for ITMX. For temporary alignment, it's OK though.

- [Sec 9-11] RIght now there is no mounts to place LO3/LO4/AS2/AS3/BHDBS. But we probably want to test something before the installation of the BHD? Just place the BHDBS on a optics mount so that we get an interfered beam on ITMY?

At this point we are supposed to have all the electronics all the CDS necessary for the new SOS control. Otherwise, they are just swinging and the alignment work will just be impossible.

- [Sec 15] The OPLEV mirrors can be freely moved as long as it does not block the main IR beams. Moving ITMXOL1 makes the reflection blocked by ITMXOL2. And moving ITMXOL2 would make the IR beams clipped. Consider replacing the mounts with a fixed mount. (The OPLEV mirrors are 1.5" in dia. It is not common vacuum compatible 1.5inch mounts. If 1" Al mirror is sufficient, we can use it.

https://wiki-40m.ligo.caltech.edu/vent/Fall2021/FinalAlignment

- The arms are the most strict alignment requirement. Everything else will follow the arm alignment. So start from the arms and propagate the alignment to Michelson / PRMI / SRMI.

- We reestablish arm alignment using the end green beams.

- Then recover IR arm alignment. Consider using ASS if possible

  16441   Sun Oct 31 14:21:31 2021 ranaHowToTreasureIFOCad

IFOcad model/video of the AEI 10m interferometer:

https://10m.aei.mpg.de/design-and-sensitivity/

  16440   Fri Oct 29 14:39:37 2021 AnchalSummaryBHD1Y1 cleared. IY3 ready for C1SUS2 I/O and FE.

[Anchal, Paco]

We cleared 1Y1 rack today removing the following items. This stuff is sitting on the floor about 2 meters east of 1Y3 (see attachment 1):

  • A VME crate: We disconnected it's power cords from the side bus.
  • A NI PXIe-1071 crate with some SMA multiplexer units on it.

We also moved the power relay ethernet strip from the middle of the rack to the bottom of the rack clearing the space marked clear in Koji's schematics. See attachment 2.

There was nothing to clear in 1Y3. It is ready for installing c1sus2 I/O chassis and FE once the testing is complete.

We also removed some orphaned hanging SMA RG-405 cables between 1Y3 and 1Y1.

Attachment 1: RemovedStuff.jpeg
RemovedStuff.jpeg
Attachment 2: 1Y1.jpeg
1Y1.jpeg
Attachment 3: 1Y3.jpeg
1Y3.jpeg
  16439   Thu Oct 28 23:13:19 2021 YehonathanUpdateBHDSOS assembly

{Tega, Yehonathan, Koji}

Suspension of a Dummy Optic

We gathered the components needed for suspending a 2" optic housed in a 3" ring adapter:

1. SOS Tower

2. Winches+Winch adapter plate (Attachment 3)

2. HeNe Laser + Power supply

3. A steering mirror

4. QPD + ND2 Filter + Electronic amplifier

5. Oscilloscope

6. Dummy optic housed in a ring adapter (Attachment 2)

An OpLev was built using the HeNe laser, steering mirror, suspended optic, QPD (Attachment 1).

The beam height was set with a ruler. The QPD was placed in front of the laser and its height was adjusted until the Y signal vanished.

The beam was made parallel to the table. First, roughly by using an iris. Then, more accurately, by measuring the beam directed at the suspended optic with the QPD and zeroing the Y signal by adjusting the steering mirror.

We suspended the dummy optic using a music wire we found in the cleanroom cabinet.

All the wire clamps on the ring adapter, SOS, and winches were loosened.

About 0.5m of music wire was cut. Then, the wire was threaded through the wire clamp on the adapter side block, around the ring adapter, and through the wire clamp on the other side block, keeping the wire untwisted. The optic was put on the safety stop and the height was roughly adjusted using the lower EQ stops.

The wire pair was then threaded through the wire clamp on the suspension block and held near the winches. The wire was made to go through the groves at the side blocks. The wire was clamped to the side blocks. The wires were clamped on the winches. The wires were pulled by the winches until the magnet marking on the side blocks was centered on the OSEMs' holes on the side plates of the SOS (attachment 4).

Once the heights were set, the wire was clamped at the suspension block.

We balanced the dummy optic using a counterweight.

We placed the QPD at the laser reflection from the optic. We adjusted the screw set going through the counterweights to adjust the balance. The pitch imbalance was monitored by observing the Y signal on the oscilloscope.

Comments:

1. Turns out that for a 1/4" thick optic, the rear counterweight needs to be removed for achieving balance.

2. To fix the counterweight on the setscrew we will use some epoxy.

3. Seems like the left optical table in the cleanroom is slightly tilted. We need to fix it.

4. The bottom long EQ stop is not touching the adapter due to the balancing mass socket. We might want to put a nut on it.

5. Still need to glue 3 more side magnet+dumbell assemblies to rods and then glue them before gluing them to the side blocks which will take at least 2 more days until we can start suspending a real optic.

Attachment 1: OpLev_Setup.jpg
OpLev_Setup.jpg
Attachment 2: suspended_optic.png
suspended_optic.png
Attachment 3: 20211029_113011.jpg
20211029_113011.jpg
Attachment 4: 20211029_113052.jpg
20211029_113052.jpg
  16438   Thu Oct 28 17:01:54 2021 AnchalSummaryBHDPart III of BHR upgrade - Adding temp fixed flat mirror for PR2

[Anchal, Paco, Ian]

  • We added a Y1-2037-0 mirror (former IPPOS SM2 mirror) on a fixed mount in the position of where PR2 is supposed to be in new BHR layout.
  • After turning out all lights in the lab, we were able to see a transmitted beam on our beam finder card.
  • We aligned the mirror so that it relfects the beam off to PR3 clearly and the reflection from PR3 hits BS in the center.
  • We were able to see clear gaussian beams splitted from BS going towards ITMX and ITMY.

Photos: https://photos.app.goo.gl/cKdbtLGa9NtkwqQ68

  16437   Thu Oct 28 16:32:32 2021 PacoSummaryBHDPart IV of BHR upgrade - Removal of BSC eastern optics

[Ian, Paco, Anchal]

We turned off the BSC oplev laser by turning the key counterclockwise. Ian then removed the following optics from the east end in the BSC:

  • OM4-PJ (wires were disconnected before removal)
  • GRX_SM1
  • OM3
  • BSOL1

We placed them in the center-front area of the XEND flow bench.

Photos: https://photos.app.goo.gl/rjZJD2zitDgxBfAdA

  16436   Wed Oct 27 19:34:52 2021 KojiSummaryElectronicsNew electronics racks

1. The rack we cleaned today (came from West Bridge) will be placed between 1X3 and 1X4, right next to 1X4 (after removing the plastic boxes). (Attachment 1)
For easier work at the side of the 1X4, the side panel of the 1X4 should be removed before placing the new rack. Note that this rack is imperial and has 10-32 threads

2. In terms of the other rack for the Y arm, we found the rack in the storage is quite dirty. Anchal pointed out that we have a few racks standing along the Y arm (as the storage of the old VME/Euro card electronics) (Attachments 2/3)
They are not too dirty and also doing nothing there. Let's vacate one of them (the one right next to the optics preparation table). Use this space as a new storage area placing a wire shelving rack for something.

BTW, I thought it is good to have the rack at the vertex side of 1Y1 (as 1Y0?), but the floor has "KEEP OUT" marking. I have no idea why we have this marking. Is this for crane operation??? Does any one know?

Attachment 1: P_20211027_180737.jpg
P_20211027_180737.jpg
Attachment 2: P_20211027_180443.jpg
P_20211027_180443.jpg
Attachment 3: P_20211027_180408.jpg
P_20211027_180408.jpg
Attachment 4: P_20211027_180139.jpg
P_20211027_180139.jpg
  16435   Wed Oct 27 18:16:45 2021 KojiSummaryBHDPart II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment

- Moving the MMT2 south by a cm is fine. This will give you ~0.5cm at TT1 without changing the other alignment much.
- IMC mode is moving because of your presence + HEPA blow.
- 2cm at Faraday is plenty for the beam diameter of a few mm.

 

  16434   Wed Oct 27 18:11:37 2021 KojiSummaryBHDPart II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment

Closed the PSL shutter @18:11
During the vent, we want to keep the cavity unlocked if not necessary.

 

  16433   Wed Oct 27 16:38:02 2021 AnchalSummaryBHDPart II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment

[Anchal, Paco]

We opened BSC and Injection Chamber doors. We removed two stacked counterweights from near the center of the BS table, from behind TT2 and placed them in the Xend flow bench. Then we unscrewed TT2 and relocated it to the new BHR layout position. This provided us with the target for the alignment of MMT1 and MMT2 mirrors.

While aligning MMT1 and MMT2, we realized that the BHR layout underestimated the clearance of the beam from MMT2 to TT2, from the TT1 suspension unit. The TT1 suspension stage was clipping our beam going to TT2. To rectify this, we decided to move the MMT2 mirror mount about a cm South and retry. We were able to align the beam to the TT2 optic, but it is a bit off-center. The reflection of TT2 now is going in the general direction of the ITMX chamber. We stopped our work here as fatigue was setting in. Following are some thoughts and future directions:

  • We realized that the output beam from the mode cleaner moves a lot (by more than a cm at MMT2) between different locks. Maybe that's just because of our presence. But we wonder how much clearance all beams must have from MC3 to TT2.
  • Currently, we think the Faraday Isolator might be less than 2 cm away from the beam between MMT1 and MMT2 and the TT1 suspension is less than 2 cm away from MMT2 and TT2.
  • Maybe we can fix these by simply changing the alignment on TT1 which was fixed for our purposes.
  • We definitely need to discuss the robustness of our path a bit more before we proceed to the next part of the upgrade.

Thu Oct 28 17:00:52 2021 After Photos: https://photos.app.goo.gl/wNL4dxPyEgYTKQFG9

  16432   Wed Oct 27 16:31:35 2021 AnchalSummaryBHDPart III of BHR upgrade - Removal of PR2 Small Suspension

I went inside the ITMX Chamber to read off specs from PR2 edge. This was required to confirm our calculations of LO power for BHR later. The numbers that I could read from the edge were kind of meaningless "0.5 088 or 2.0 088". To make it more worthwhile this opening of the chamber, we decided to remove the PR2 suspension unit so that the optic can be removed and installed on an SOS in the cleanroom. We covered the optic in clean aluminum foil inside the chamber, then placed in on another aluminum foil to cover completely. Then I traveled slowly to the C&B room, where I placed it on a flow bench.


Later on, we decided to use a dummy fixed mount mirror for PR2 initially with the same substrate thickness, so that we get enough LO power in transmission for alignment. In the very end, we'll swap that with the PR2 mounted on an SOS unit.

  16431   Wed Oct 27 16:27:16 2021 AnchalSummaryBHDPart II of BHR upgrade - Prep

[Anchal, Paco, Ian]

Before we could start working on Part II, which is to relocate TT2 to new location, we had to clear space in front of injection chamber door and clean the floor which was very dusty. This required us to disconnect everything we could safely from OMC North short electronics rack, remove 10-15 BNC cables, 4-5 power cords and relocate some fiber optic cables. We didn't had caps for fiber optic cables handy, so we did not remove them from the rack mounted unit and just turned it away. At the end, we mopped the floor and dried it with a dry cloth. Before and after photos in attachments.

 

Attachment 1: OMCNorthBefore.jpeg
OMCNorthBefore.jpeg
Attachment 2: OMCNorthAfter.jpeg
OMCNorthAfter.jpeg
  16430   Tue Oct 26 18:24:00 2021 Ian MacMillanSummaryCDSc1sus2 DAC to ADC test

[Ian, Anchal, Paco]

After the Koji found that there was a problem with the power source Anchal and I fixed the power then reran the measurment. The only change this time around is that I increased the excitation amplitude to 100. In the first run the excitation amplitude was 1 which seemed to come out noise free but is too low to give a reliable value.

link to previous results

The new plots are attached.

Attachment 1: data2_Plots.pdf
data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf
Attachment 2: data3_Plots.pdf
data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf
  16429   Tue Oct 26 16:56:22 2021 PacoSummaryBHDPart I of BHR upgrade - Locked PMC and IMC

[Paco, Ian]

We opened the laser head shutter. Then, we scanned around the PMC resonance and locked it. We then opened the PSL shutter, touched the MC1, MC2 and MC3 alignment (mostly yaw) and managed to lock the IMC. The transmission peaked at ~ 1070 counts (typical is 14000 counts, so at 10% of PSL power we would expect a peak transmission of 1400 counts, so there might still be some room for improvement). The lock was engaged at ~ 16:53, we'll see for how long it lasts.

There should be IR light entering the BSC!!! Be alert and wear laser safety goggles when working there.

We should be ready to move forward into the TT2 + PR3 alignment.

  16428   Tue Oct 26 14:53:24 2021 KojiUpdateElectronicsRack

1. We have a rack at the 40m storage. We are free to take it to the lab. If there is a tag, tell the info to Liz. Let's move it to the lab tomorrow right after the meeting.

2. We have a few racks in WB B1 (Attachment 1). Liz and I checked a rack which looks suitable for us. 46U height. Caltech transport will move it to the lab.

Attachment 1: P_20211026_143814.jpg
P_20211026_143814.jpg
  16427   Tue Oct 26 13:27:07 2021 TegaSummaryElectronicsSat Amp modification Summary

Modifications and testing of SatAmp units COMPLETE. Attachments 1 & 2 show all 19 units, one installed unit and the remaining 18 units are stacked and ready for install. Detailed notes of the modification for each unit are presented in the summary document in the dcc.

 

 

Attachment 1: SapAmpModStack.jpg
SapAmpModStack.jpg
Attachment 2: SatAmpInstalled.jpg
SatAmpInstalled.jpg
  16426   Tue Oct 26 10:17:14 2021 YehonathanUpdateBHDSOS assembly

Things that I need to start suspending optics:

1. Winch adapter plate (D970314). Might need to make one.

2. Quad photodetector

3. Camera and camera mount

4. Beam height target

5. Height gauge

  16425   Mon Oct 25 17:37:42 2021 AnchalSummaryBHDPart I of BHR upgrade - Removed optics from BSC

[Anchal, Paco, Ian]


Clean room etiquettes

  • Two people in coverall suits, head covers, masks and AccuTech ultra clean gloves.
  • One person in just booties to interact with outside "dirty" world.
  • Anything that comes in chamber, first cleaned outside with clean cloth and IPA. Then cleaned by the "clean" folks. We followed this for allen keys, camera and beam finder card.
  • Once the chamber cover has been removed, cover the annulus with donut. We forgot to do this :(

Optics removal and changes

We removed the following optics from the BSC table and stored them in X-end flowbench with fan on. See attachment 1 and 2.

  1. IPPOS SM2
  2. GRX SM2
  3. PRM OL1
  4. PRMOL4
  5. IPPOS SM3
  6. IPANG SM1
  7. PRM OL2
  8. Unidentified optic inbetween IPPOS45P and IPPOS SM3
  9. Beam block behing PR3
  10. Beam block behind GR PBS
  11. GR PBS
  12. GRPERI1L (Periscope)
  13. PRMOL3
  14. IPPOS45P
  15. Cylindrical counterweight on North-west end of table.
  16. Cheap rectangular mirror on South west end of table (probably used for some camera, but not in use anymore)
  17. IPANGSM2

We also changed the direction of clamp of MMT1 to move it away from the center of the able (where PRM will be placed)

We screwed in the earthquake stops on PRM and BS from front face and top.

We unscrewed the cable post for BS and PRM oplevs and loved it in between SR3 and BS and screwed it lightly.

We moved the PRM, turned it anti-clockwise 90 degrees and brought it in between TT2 and BS. Now there is a clear line of sight between TT2 and PR2 on ITMY table.


Some next steps:

  • We align the input beam to TT2 by opening the "Injection Chamber" (formerly known as OMC chamber). While doing so, we'll clear unwanted optics from this table as well.
  • We open ITMX chamber, clear some POP optics. If SOS are ready, we would replace PR2 with SOS and put it in a new position.
  • Then we'll replace PR3 with an SOS and align the beam to BS.

These are next few days of work. We need atleast one SOS ready by Thursday.


Photos after today's work: https://photos.app.goo.gl/EE7Mvhw5CjgZrQpG6

Attachment 1: rn_image_picker_lib_temp_44cb790a-c3b4-42aa-8907-2f9787a02acd.jpg
rn_image_picker_lib_temp_44cb790a-c3b4-42aa-8907-2f9787a02acd.jpg
Attachment 2: rn_image_picker_lib_temp_0fd8f4fd-64ae-4ccd-8422-cfe929d4eeee.jpg
rn_image_picker_lib_temp_0fd8f4fd-64ae-4ccd-8422-cfe929d4eeee.jpg
  16424   Mon Oct 25 13:23:45 2021 AnchalSummaryBHDBefore photos of BSC

[Yehonathan, Anchal]

On thursday Oct 21 2021, Yehonathan and I opened the door to BSC and took some photos. We setup the HEPA stand next to the door with anti-static curtains covering all sides. We spend about 15 minutes trying to understand the current layout and taking photos and a video. Any suggestions on improvement in our technique and approach would be helpful.

Links to photos:

https://photos.app.goo.gl/fkkdu9qAvH1g5boq6

  16423   Fri Oct 22 17:35:08 2021 Ian MacMillanSummaryPEMParticle counter setup near BS Chamber

I have done some reading about where would be the best place to put the particle counter. The ISO standard (14644-1:2015) for cleanrooms is one every 1000 m^2 so one for every 30m x 30m space. We should have the particle counter reasonably close to the open chamber and all the manufactures that I read about suggest a little more than 1 every 30x30m. We will have it much closer than this so it is nice to know that it should still get a good reading. They also suggest keeping it in the open and not tucked away which is a little obvious. I think the best spot is attached to the cable tray that is right above the door to the control room. This should put it out of the way and within about 5m of where we are working. I ordered some cables to route it over there last night so when they come in I can put it up there. 

  16422   Thu Oct 21 15:24:35 2021 ranaSummaryPEMParticle counter setup near BS Chamber

rethinking what I said on Wednesday - its not a good idea to put the particle counter on a vac chamber with optics inside. The rumble from the air pump shows up in the acoustic noise of the interferometer. Let's look for a way to mount it near the BS chamber, but attached to something other than vacuum chambers and optical tables.

Quote:

I have placed a GT321 particle counter on top of the MC1/MC3 chamber next to the BS chamber.

 

  16421   Thu Oct 21 15:22:35 2021 ranaSummaryPEMParticle counter setup near BS Chamber

SVG doesn't work in my browser(s). Can we use PDF as our standard for all graphics other than photos (PNG/JPG) ?

  16420   Thu Oct 21 11:41:31 2021 AnchalSummaryPEMParticle counter setup near BS Chamber

The particle count channel names were changes yesterday to follow naming conventions used at the sites. Following are the new names:

C1:PEM-BS_DUST_300NM
C1:PEM-BS_DUST_500NM
C1:PEM-BS_DUST_1000NM
C1:PEM-BS_DUST_2000NM
C1:PEM-BS_DUST_5000NM
 

The legacy count channels are kept alive with C1:PEM-count_full copying C1:PEM-BS_DUST_1000NM channel and C1:PEM-count_half copying C1:PEM-BS_DUST_500NM channel.

Attachment one is the particle counter trend since 8:30 am morning today when the HVAC wokr started. Seems like there was some peak particle presence around 11 am. The particle counter even counted 8 counts of particles size above 5um!

 

Attachment 1: ParticleCountData20211021.pdf
ParticleCountData20211021.pdf
  16419   Thu Oct 21 11:38:43 2021 JordanUpdateSUSStandoffs for Side Magnet on 3" Adapter Ring SOS Assembly

I had 8 standoffs made at the Caltech chemistry machine shop to be used as spacers for the side magnets on the 3" Ring assembly. This is to create enough clearance between the magnet and the cap screws directly above on the wire clamp.

These are 0.075" diameter by .10" length. Putting them through clean and bake now.

Attachment 1: Magnet_Standoffs.jpg
Magnet_Standoffs.jpg
  16418   Wed Oct 20 15:58:27 2021 KojiUpdateVACHow to vent TP1

Probably the hard disk of c0rga is dead. I'll follow up in this elog later today.

Looking at the log in /opt/rtcds/caltech/c1/scripts/RGA/logs , it seemed that the last RGA scan was Sept 2, 2021, the day when we had the disk full issue of chiara.
I could not login to c0rga from control machines.
I was not aware of the presence for c0rga until today, but I could locate it in the X arm.
The machine was not responding and it was rebooted, but could not restart. It made some knocking sound. I am afraid that the HDD failed.

I think we can
- prepare a replacement linux machine for the python scripts
or
- integrate it with c1vac

  16417   Wed Oct 20 11:48:27 2021 AnchalSummaryCDSPower supple configured correctly.

This was horrible! That's my bad, I should have checked the configuration before assuming that it is right.

I fixed the power supply configuration. Now the strip has two rails of +/- 18V and the GND is referenced to power supply earth GND.

Ian should redo the tests.

  16416   Wed Oct 20 11:16:21 2021 AnchalSummaryPEMParticle counter setup near BS Chamber

I have placed a GT321 particle counter on top of the MC1/MC3 chamber next to the BS chamber. The serial cable is connected to c1psl computer on 1X2 using 2 usb extenders (blue in color) over the PSL enclosure and over the 1X1 rack.

The main serial communication script for this counter by Radhika is present in 40m/labutils/serial_com/gt321.py.

A 40m specific application script is present in the new git repo for 40m scripts, in 40m/scripts/PEM/particleCounter.py. Our plan is to slowly migrate the legacy scripts directory to this repo overtime. I've cloned this repo in the nfs shared directory at /opt/rtcds/caltech/c1/Git/40m/scripts which makes the scripts available at all computers and keep them upto date in all computers.

The particle counter script is running on c1psl through a systemd service, using service file 40m/scripts/PEM/particleCounter.service. Locally in c1psl, /etc/systemd/system/particleCounter.service is symbollically linked to the file in the file.

Following channels for particle counter needed to be created as I could not find any existing particle counter channels.

[C1:PEM-BS_PAR_CTS_0p3_UM]
[C1:PEM-BS_PAR_CTS_0p5_UM]
[C1:PEM-BS_PAR_CTS_1_UM]
[C1:PEM-BS_PAR_CTS_2_UM]
[C1:PEM-BS_PAR_CTS_5_UM]

These are created from 40m/softChansModbus/particleCountChans.db database file. Computer optimus is running a docker container to serve as EPICS server for such soft channels. To add or edit channels, one just need to add new database file or edit database files in thsi repo and on optimus do:

controls@optimus|~> sudo docker container restart softchansmodbus_SoftChans_1
softchansmodbus_SoftChans_1

that's it.

I've added the above channels to /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini to record them in framebuilder. Starting from 11:20 am Oct 20, 2021 PDT, the data on these channels is from BS chamber area. Currently the script is running continuosly, which means 0.3u particles are sampled every minute, 0.5u twice in 5 minutes and 1u, 2u, and 5u particles are sampled once in 5 minutes. We can reduce the sampling rate if this seems unncessary to us.

Attachment 1: PXL_20211020_183728734.jpg
PXL_20211020_183728734.jpg
  16415   Tue Oct 19 23:43:09 2021 KojiSummaryCDSc1sus2 DAC to ADC test

(Because of a totally unrelated reason) I was checking the electronics units for the upgrade. And I realized that the electronics units at the test stand have not been properly powered.

I found that the AA/AI stack at the test stand (Attachment 1) has an unusual powering configuration (Attachment 2).
- Only the positive power supply was used / - The supply voltage is only +15V / - The GND reference is not connected to anywhere.

For confirmation, I checked the voltage across the DC power strip (Attachments 3/4). The positive was +5.3V and the negative was -9.4V. This is subject to change depending on the earth potential.

This is not a good condition at all. The asymmetric powering of the circuit may cause damages to the opamps. So I turned off the switches of the units.

The power configuration should be immediately corrected.

  1. Use both positive and negative supply (2 power supply channels) to produce the positive and the negative voltage potentials. Connect the reference potential to the earth post of the power supply.
    https://www.youtube.com/watch?v=9_6ecyf6K40   [Dual Power Supply Connection / Serial plus minus electronics laboratory PS with center tap]
  2. These units have DC power regulator which produces +/-15V out of +/-18V. So the DC power supplies are supposed to be set at +18V.

 

Attachment 1: P_20211019_224433.jpg
P_20211019_224433.jpg
Attachment 2: P_20211019_224122.jpg
P_20211019_224122.jpg
Attachment 3: P_20211019_224400.jpg
P_20211019_224400.jpg
Attachment 4: P_20211019_224411.jpg
P_20211019_224411.jpg
  16414   Tue Oct 19 18:20:33 2021 Ian MacMillanSummaryCDSc1sus2 DAC to ADC test

I ran a DAC to ADC test on c1sus2 channels where I hooked up the outputs on the DAC to the input channels on the ADC. We used different combinations of ADCs and DACs to make sure that there were no errors that cancel each other out in the end. I took a transfer function across these channel combinations to reproduce figure 1 in T2000188.

As seen in the two attached PDFs the channels seem to be working properly they have a flat response with a gain of 0.5 (-6 dB). This is the response that is expected and is the result of the DAC signal being sent as a single ended signal and the ADC receiving as a differential input signal. This should result in a recorded signal of 0.5 the amplitude of the actual output signal.

The drop off on the high frequency end is the result of the anti-aliasing filter and the anti-imaging filter. Both of these are 8-pole elliptical filters so when combined we should get a drop off of 320dB per decade. I measured the slope on the last few points of each filter and the averaged value was around 347dB per decade. This is slightly steeper than expected but since it is to cut off higher frequencies it shouldn't have an effect on the operation of the system. Also it is very close to the expected value.

The ripples seen before the drop off are also an effect of the elliptical filters and are seen in T2000188.

Note: the transfer function that doesn't seem to match the others is the heartbeat timing signal.

Attachment 1: data3_Plots.pdf
data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf data3_Plots.pdf
Attachment 2: data2_Plots.pdf
data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf data2_Plots.pdf
  16413   Tue Oct 19 11:30:39 2021 KojiUpdateVACHow to vent TP1

I learned that TP1 was vented through the RGA room in the past. This can be done by opening VM2 and a manual valve ("needle valve")
I checked the setup and realized that this will vent RGA. But it is OK as long as we turns of the RGA during vent and bake it once TP1 is back.

Additional note:

- It'd be nice to take a scan for the current background level before the work.
- Turn RGA EM and filament off, let it cool down overnight. 
- Vent with clean N2 or clean air. (Normal operating temp ~80C is to minimize accumulation of H-C contaminations.)
- There is a manual switch and indicators on the top of the RGA amp. It has auto protection to turn filament off if the pressure increase over ~1e-5.

Attachment 1: Screen_Shot_2021-10-18_at_14.52.34.png
Screen_Shot_2021-10-18_at_14.52.34.png
ELOG V3.1.3-